OpenStudy (ksaimouli):

determine the exact value? sec(-5pie/4)

OpenStudy (ksaimouli):

\[\sec(-5\Pi/4)\]

OpenStudy (jamesj):

sec(x) = 1/cos(x) and cos(pi+x) = -cos(x) Hence cos(5pi/4) = -cos(pi/4) = -1/sqrt(2) and sec(5pi/4) = -sqrt(2)